please tell me the formulas for solvilng the question related to binomial coefficient?

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

Here are the formulas for binomial coefficient:
1)nCr= [n!/(r!(n-r)!)]
2) nCr= nCs
Either r = s or r + s = n
3) nCr+ nCr+1= n+1Cr+1
4)n+1Cr+1= (n+1/r+1)nCr
5) nCr+1/nCr= n-r/r+1